An oil prospecting firm hits oil or gas on 10% of its drillings. If the firm drills two wells. the four possible sim accompanying table. Find the probability that the company will hit oil or gas 2.15 ple events and three of their associated probabilities are as given in the a on the first drilling and miss on the secon<d b on at least one of the two drillings. Outcome of SimpleOutcome of Event First Drilling Second Drilling Probability 01 E,Hit (oil or gas) E2Hit Hit (oil or gas) ISS Miss Miss 09 81 it E. Miss
Answer #1

a)P(hit on first drill and miss on second drill) =P(E2) =1-0.01-0.09-0.81=0.10

b)P(on at least one of two drillings) =1-E(4)=1-0.81 =0.19